OpenStudy (anonymous):

solve the equation. 3x-7=5x+19. check your answer.

OpenStudy (anonymous):

I don't understand

OpenStudy (lazyboy):

forget that then.. here is a more detailed explanation..

OpenStudy (anonymous):

ok

OpenStudy (lazyboy):

Simplifying 3x + -7 = 5x + 19 Reorder the terms: -7 + 3x = 5x + 19 Reorder the terms: -7 + 3x = 19 + 5x Solving -7 + 3x = 19 + 5x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5x' to each side of the equation. -7 + 3x + -5x = 19 + 5x + -5x Combine like terms: 3x + -5x = -2x -7 + -2x = 19 + 5x + -5x Combine like terms: 5x + -5x = 0 -7 + -2x = 19 + 0 -7 + -2x = 19 Add '7' to each side of the equation. -7 + 7 + -2x = 19 + 7 Combine like terms: -7 + 7 = 0 0 + -2x = 19 + 7 -2x = 19 + 7 Combine like terms: 19 + 7 = 26 -2x = 26 Divide each side by '-2'. x = -13 Simplifying x = -13 If this helped click on best response

OpenStudy (anonymous):

i am so not good in math sry

OpenStudy (lazyboy):

did u get it though?

OpenStudy (anonymous):

hold on i am reading everything so i can understand what u got

OpenStudy (lazyboy):

ok

OpenStudy (lazyboy):

so?..

OpenStudy (anonymous):

hold on i am trying to solve it as wll so i under stand all of it

OpenStudy (lazyboy):

alright..

OpenStudy (anonymous):

ok i got it but how do i check my answer

OpenStudy (lazyboy):

hmm

OpenStudy (lazyboy):

I'm not sure, I am only good at simplifying but not checking. I never do it sry :/

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do you know how to check the work?